find domain and range of 1/(root of x3-4x)
Answer by stanbon(75887)

find domain and range of 1/(root of x^3-4x)
------
x^3-4x >= 0
x(x^2-4) >= 0
x(x-2)(x+2) >= 0
----
Draw a number line.
Plot x = -2; x = 0; x = 2
----
Test 4 value intervals on the line:
x(x-2)(x+2) >= 0
x = -10, you get -*-*- > 0 ; false
x = -1, you get -*-*+ >0 ; true
x = +1, you get -*+*+ >0 0 ; false
x = 10, you get +*+*+ > 0 ; true
----
Domain: All Real Numbers [-2,0] U {2.+oo)
Range: All Real Numbers >= 0
====================================
========================================
Cheers,
Stan H.